# Efficient designs for multivariate crossover trials

### v1

@@ 2 A multivariate crossover model
Here, we consider $g=2$. Let $\sigma_{11}$ and $\sigma_{22}$, where $\sigma_{11},\sigma_{22}>0$, be the respective variances of the first and second responses and $\rho$, $10<|\rho|<$ 1, denotes the correlation coefficient between the observations from the two responses measured from the same period. Under the generalized markov-type structure the dispersion matrix $\bm{\varSigma}$ is as follows:
$\bm{\varSigma}$ $=\begin{bmatrix}\bm{\varSigma}_{11}&\bm{\varSigma}_{12}\\ \bm{\varSigma}_{21}&\bm{\varSigma}_{22}\end{bmatrix},$
where $\bm{\varSigma}_{11}=\bm{I}_{n}\otimes\bm{V}_{1}$, $\bm{\varSigma}_{12}=\bm{\varSigma}_{21}=\rho\sqrt{\frac{\sigma_{22}}{\sigma_{11}}}\bm{\varSigma}_{11}$, $\bm{\varSigma}_{22}=\rho^{2}\frac{\sigma_{22}}{\sigma_{11}}\bm{\varSigma}_{11}+\sigma_{22}(1-\rho^{2})\bm{\varSigma}_{R}$ and $\bm{\varSigma}_{R}=\bm{I}_{n}\otimes\bm{V}_{R}$. Here, $\bm{V}_{1}=\sigma_{11}\bm{V}_{C}$, $\bm{V}_{C}$ is a known positive definite and symmetric matrix, and $\bm{V}_{R}$ is a known correlation matrix. We see that in this structure, direct covariances for both responses may differ, while the cross covariances are proportional to the direct covariance within the first response. Usually in cokriging spatial experiments the first response is taken to be the primary response (Journel, 1999; Chiles and Delfiner, 2012; Dasgupta et al., 2022). In our case both responses are given the same importance.
@@ 3 Search for efficient designs
We next determine a trace optimal/efficient design for the direct effects in model (2.4). Using the idea as provided by Bate and Jones, (2006), we say that a design $d^{*}\in\mathcal{D}$, where $\mathcal{D}$ is a subclass of designs, is a trace optimal/efficient design for the parameters of interest over $\mathcal{D}$, if $d^{*}$ maximizes the trace of the corresponding information matrix over $\mathcal{D}$.
For any design $d\in\varOmega_{t,n,p}$, let $\bm{C}_{d(s1)}$ and $\bm{C}_{d(s2)}$ represent the information matrices of the direct effects for the proportional and generalized markov-type structures, respectively. Lemmas S.1 and S.3 show that the two information matrices have the following representations:
@@
$\bm{C}^{{}^{\prime}}_{{d(s2)}(21)(1)}=\left(\bm{I}_{2}\otimes\bm{T}_{d}^{{}^{\prime}}\right)\bm{A}^{*}\left(\bm{I}_{2}\otimes\bm{F}_{d}\bm{H}_{t}\right)$, $\bm{C}_{{d(s2)}(22)(1)}=\left(\bm{I}_{2}\otimes\bm{F}^{{}^{\prime}}_{d}\bm{H}_{t}\right)\bm{A}^{*}\left(\bm{I}_{2}\otimes\bm{F}_{d}\bm{H}_{t}\right)$, and the matrix $\bm{A}^{*}$ is $\bm{\varSigma}^{-1/2}pr^{\perp}\left(\bm{\varSigma}^{-1/2}\bm{Z}_{1}\right)\bm{\varSigma}^{-1/2}$. Here, $\bm{H}_{t}=\bm{I}_{t}-\frac{1}{t}\bm{{1}}_{t}\bm{{1}}{t}^{{}^{\prime}}$. Note that $\bm{A}^{*}$ depends on the structure of $\bm{\varSigma}$ and $\bm{C}_{d(s1)}$ is a $gt\times gt$ matrix, while $\bm{C}_{d(s2)}$ is a $2t\times 2t$ matrix.
@@ 3.1 Proportional covariance structure
$\bm{A}^{*}$ $=\bm{\varGamma}^{-1}\otimes\left(\bm{H}_{n}\otimes\bm{V}^{*}\right),$
@@
$\bm{C}_{d(uni)(12)}=\bm{C}_{d(uni)(21)}^{{}^{\prime}}=\bm{T}_{d}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{V}^{*}\right)\bm{F}_{d}$, and $\bm{C}_{d(uni)(22)}=\bm{F}_{d}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{V}^{*}\right)\bm{F}_{d}$. Here, $\bm{V}^{*}=\bm{V}^{-1}-\left(\bm{{1}}_{p}^{{}^{\prime}}\bm{V}^{-1}\bm{{1}}_{p}\right)^{-1}\bm{V}^{-1}\bm{J}_{p\times p}\bm{V}^{-1}$.
@@
It should be noted $\bm{C}_{d(uni)}$ is the information matrix for the direct effects when we have a univariate version of the model (2.1), i.e., $g=1$. To check if $OA_{I}\left(n=\lambda t\left(t-1\right),p=t,t,2\right)$, where $\lambda$ is a positive integer and $t\geq 3$, is universally optimal, we explore complete symmetricity of the corresponding information matrix for $g>1$ case.
@@
where $\gamma^{(kk^{\prime})}$ is the $(k,k^{\prime})^{th}$ element of the matrix $\bm{\varGamma}^{-1}$. However, this condition is also seen to be violated for the $g>1$ setup (see Proof of Remark 3.1). Note, for a multivariate setup with no correlation between responses and a much simpler covariance structure, Niphadkar and Mukhopadhyay, (2023) have successfully obtained universally optimal designs. However, this is not true when the responses are correlated. Hence, for the multivariate response case, we resort to identifying an efficient design in the next theorem.
@@
$\mathcal{D}^{(1)}_{t,n=\lambda t(t-1),p=t}$ is a class of binary designs with $p=t$, $\lambda$ is a positive integer and $t\geq 3$. It has been shown by Kunert and Martin, (2000) that $d^{*}$ is universally optimal for the direct effects over $\mathcal{D}^{(1)}_{t,n=\lambda t(t-1),p=t}$ in the univariate response case. Note that universal optimality of a design over a subclass of designs implies that the design maximizes the trace of the corresponding information matrix over the same subclass of designs. Hence, $d^{*}$ maximizes $tr\left(\bm{C}_{d(uni)}\right)$ over $\mathcal{D}^{(1)}_{t,n=\lambda t(t-1),p=t}$, where $\bm{C}_{d(uni)}$ is the information matrix of the direct effects for the univariate response case . For $g>1$, using the expression of $\bm{C}_{d(s1)}$ from Theorem 3.1, we get
$tr\left(\bm{C}_{d(s1)}\right)$ $=tr\left(\bm{\varGamma}^{-1}\right)tr\left(\bm{C}_{d(uni)}\right).$ (3.4)
Here, $\bm{\varGamma}^{-1}$ is a positive definite matrix, and thus $tr\left(\bm{\varGamma}^{-1}\right)>0$. So, from (3.4), we get that for $g>1$, $d^{*}$ maximizes $tr\left(\bm{C}_{d(s1)}\right)$ over $\mathcal{D}^{(1)}_{t,n=\lambda t(t-1),p=t}$. Thus $d^{*}$ is trace optimal/efficient for the direct effects over $\mathcal{D}^{(1)}_{t,n=\lambda t(t-1),p=t}$ for the $g>1$ case. ∎
@@ 3.2 Generalized markov-type covariance structure
For the markovian structure, we are able to show that the $OA_{I}\left(n=\lambda t\left(t-1\right),p=t,t,2\right)$, where $\lambda$ is a positive integer and $t\geq 3$, is a highly efficient design. To evaluate the trace optimality/efficiency of these $OA_{I}\left(n=\lambda t\left(t-1\right),p=t,t,2\right)$ designs, we obtain an upper bound of $tr\left(\bm{C}_{d(s2)}\right)$, and state a necessary and sufficient condition for a $OA_{I}\left(n=\lambda t\left(t-1\right),p=t,t,2\right)$, to attain this upper bound of $tr\left(\bm{C}_{d(s2)}\right)$. Note, that if this upper bound is attained then the design maximizes $tr\left(\bm{C}_{d(s2)}\right)$ and hence is an efficient design. Also, for various choices $\bm{V}_{1}$, $\bm{V}_{R}$ and $\rho$, we study the trace of the information matrix for the $OA_{I}\left(n=\lambda t\left(t-1\right),p=t,t,2\right)$ and the derived upper bound.
Proceeding as before, we see that the process of obtaining $\bm{A}^{*}$ is more complicated as compared to the proportional structure. This is mainly because under the Markovian structure, $\bm{\varSigma}_{22}$ is not proportional to $\bm{\varSigma}_{11}$. By laborious calculation, (see Lemma S.4), we obtain the expression of the matrix $\bm{A}^{*}$ under the generalized markov-type structure as
$\bm{A}^{*}$ $=\begin{bmatrix}\bm{H}_{n}\otimes\bm{\varOmega}_{1}&-\bm{H}_{n}\otimes\bm{\varOmega}_{2}\\ -\bm{H}_{n}\otimes\bm{\varOmega}_{2}&\bm{H}_{n}\otimes\bm{\varOmega}_{4}\end{bmatrix},$
where $\bm{\varOmega}_{1}=\bm{V}_{1}^{*}+\frac{\bar{\rho}^{2}}{\sigma_{12}}\bm{V}_{R}^{*}$, $\bm{\varOmega}_{2}=\frac{\bar{\rho}}{\sigma_{12}}\bm{V}_{R}^{*}$, and $\bm{\varOmega}_{4}=\frac{1}{\sigma_{12}}\bm{V}_{R}^{*}$. Here, $\bm{V}_{1}^{*}=\bm{V}_{1}^{-1}-$
$\left(\bm{{1}}_{p}^{{}^{\prime}}\bm{V}_{1}^{-1}\bm{{1}}_{p}\right)^{-1}\bm{V}_{1}^{-1}\bm{J}_{p\times p}\bm{V}_{1}^{-1}$, $\bm{V}_{R}^{*}=\bm{V}_{R}^{-1}-\left(\bm{{1}}_{p}^{{}^{\prime}}\bm{V}_{R}^{-1}\bm{{1}}_{p}\right)^{-1}\bm{V}_{R}^{-1}\bm{J}_{p\times p}\bm{V}_{R}^{-1}$, $\bm{H}_{n}=\bm{I}_{n}-\frac{1}{n}\bm{{1}}_{n}\bm{{1}}_{n}^{{}^{\prime}}$, $\bm{J}_{p\times p}=\bm{{1}}_{p}\bm{{1}}_{p}^{{}^{\prime}}$, $\bar{\rho}=\rho\sqrt{\frac{\sigma_{22}}{\sigma_{11}}}$ and $\sigma_{12}=\sigma_{22}\left(1-\rho^{2}\right)$. Note that further we use $\bm{H}_{p}=\bm{I}_{p}-\frac{1}{p}\bm{{1}}_{p}\bm{{1}}_{p}^{{}^{\prime}}$.

### v2

@@ 2 A multivariate crossover model
In the statistical literature (see Journel, 1999; Chiles and Delfiner, 2012; Dasgupta et al., 2022), this structure is available to model the $g=2$ case. Let $\sigma_{11}$ and $\sigma_{22}$, where $\sigma_{11},\sigma_{22}>0$, be the respective variances of the first and second responses and $\rho$, $10<|\rho|<$ 1, denotes the correlation coefficient between the observations from the two responses measured from the same period. Under the generalized Markov-type structure the dispersion matrix $\bm{\mathit{\Sigma}}$ is as follows:
$\bm{\mathit{\Sigma}}$ $=\begin{bmatrix}\bm{\mathit{\Sigma}}_{11}&\bm{\mathit{\Sigma}}_{12}\\ \bm{\mathit{\Sigma}}_{21}&\bm{\mathit{\Sigma}}_{22}\end{bmatrix},$
where $\bm{\mathit{\Sigma}}_{11}=\bm{I}_{n}\otimes\bm{V}_{1}$, $\bm{\mathit{\Sigma}}_{12}=\bm{\mathit{\Sigma}}_{21}=\rho\sqrt{\frac{\sigma_{22}}{\sigma_{11}}}\bm{\mathit{\Sigma}}_{11}$, $\bm{\mathit{\Sigma}}_{22}=\rho^{2}\frac{\sigma_{22}}{\sigma_{11}}\bm{\mathit{\Sigma}}_{11}+\sigma_{22}(1-\rho^{2})\bm{\mathit{\Sigma}}_{R}$ and $\bm{\mathit{\Sigma}}_{R}=\bm{I}_{n}\otimes\bm{V}_{R}$. Here, $\bm{V}_{1}=\sigma_{11}\bm{V}_{C}$, $\bm{V}_{C}$ is a known positive definite and symmetric matrix, and $\bm{V}_{R}$ is a known correlation matrix.
2.1 Working covariance structures
The two working covariance structures for the dispersion matrix $\bm{\mathit{\Sigma}}$, namely, the proportional structure and the generalized Markov-type structure are not at all restrictive, as they allow usage of various covariance functions as shown next. For the proportional case, the only constraint on the matrix $\bm{V}$ is positive definiteness and symmetricity. Some popular choices of $\bm{V}$ are shown in Table 1.
[b] Covariance function for $\bm{V}$ $\left(\bm{V}\right)_{i_{1},i_{2}}:\left(i_{1},i_{2}\right)^{th}\text{ element of }\bm{V}$ $10<r_{(1)}<$ 1, and $i_{1}$ and $i_{2}$ are period indices, $1\leq i_{1},i_{2}\leq p$. Mat($0.5$) 1,2 $r_{(1)}^{|i_{1}-i_{2}|}$ Mat($1.5$) 1 $\left[1-|i_{1}-i_{2}|log\left(r_{(1)}\right)\right]r_{(1)}^{|i_{1}-i_{2}|}$ Mat($\infty$) 1,2 $r_{(1)}^{|i_{1}-i_{2}|^{2}}$
Table 1: Some popular choices for $\bm{V}$ in the proportional covariance structure
  • 1
    See Dasgupta et al., (2022) and Li and Zimmerman, (2015).
  • 2
    Note that $r_{(1)}$ is the within and between response correlation coefficient between consecutive periods.
For the generalized Markov-type covariance, we need to choose matrices $\bm{V}_{1}$ and $\bm{V}_{R}$. Note, $\bm{V}_{1}=\sigma_{11}\bm{V}_{C}$, where $\bm{V}_{C}$ is a known positive definite and symmetric matrix, $\sigma_{11}>0$ and $\bm{V}_{R}$ is a known correlation matrix. Various popular combinations of the matrices $\bm{V}_{1}$ and $\bm{V}_{R}$ are presented in Table 2. Note that we could also use the same covariance functions for $\bm{V}_{1}$ and $\bm{V}_{R}$ with different correlation parameters between consecutive periods.
[b] Covariance function for $\bm{V}_{1}$ and $\bm{V}_{R}$ $\left(\bm{V}_{1}\right)_{i_{1},i_{2}}$: $\left(i_{1},i_{2}\right)^{th}$ element of $\bm{V}_{1}$ $\left(\bm{V}_{R}\right)_{i_{1},i_{2}}$: $\left(i_{1},i_{2}\right)^{th}$ element of $\bm{V}_{R}$ $10<r_{(1)}<$ 1, and $i_{1}$ and $i_{2}$ are period indices, $1\leq i_{1},i_{2}\leq p$. Note that $\sigma_{11}>0$ for all the structures below. (Mat($0.5$), Mat($1.5$)) $\sigma_{11}r_{(1)}^{|i_{1}-i_{2}|}$ $\left[1-|i_{1}-i_{2}|log\left(r_{(1)}\right)\right]r_{(1)}^{|i_{1}-i_{2}|}$ (Mat($0.5$), Mat($\infty$)) $\sigma_{11}r_{(1)}^{|i_{1}-i_{2}|}$ $r_{(1)}^{|i_{1}-i_{2}|^{2}}$ (Mat($1.5$), Mat($0.5$)) $\sigma_{11}\left[1-|i_{1}-i_{2}|log\left(r_{(1)}\right)\right]r_{(1)}^{|i_{1}-i_{2}|}$ $r_{(1)}^{|i_{1}-i_{2}|}$ (Mat($1.5$), Mat($\infty$)) $\sigma_{11}\left[1-|i_{1}-i_{2}|log\left(r_{(1)}\right)\right]r_{(1)}^{|i_{1}-i_{2}|}$ $r_{(1)}^{|i_{1}-i_{2}|^{2}}$ (Mat($\infty$), Mat($0.5$)) $\sigma_{11}r_{(1)}^{|i_{1}-i_{2}|^{2}}$ $r_{(1)}^{|i_{1}-i_{2}|}$ (Mat($\infty$), Mat($1.5$)) $\sigma_{11}r_{(1)}^{|i_{1}-i_{2}|^{2}}$ $\left[1-|i_{1}-i_{2}|log\left(r_{(1)}\right)\right]r_{(1)}^{|i_{1}-i_{2}|}$ (Mat($0.5$), Mat($0.5$)) 3 $\sigma_{11}r_{(1)}^{|i_{1}-i_{2}|}$ $r_{(1)}^{2|i_{1}-i_{2}|}$
Table 2: Some popular combinations of $\bm{V}_{1}$ and $\bm{V}_{R}$ under the generalized Markov-type structure
  • 3
    $\bm{V}_{1}$ and $\bm{V}_{R}$ have Mat($0.5$) structure with correlation between adjacent periods as $r_{(1)}$ and $r_{(1)}^{2}$, respectively. This structure has been named as NS1 in Li and Zimmerman, (2015) and Dasgupta et al., (2022).
To aid the reader, we use an illustration based on a crossover design with $t=p=n=3$ and $g=2$. In this case we write $\bm{\mathit{\Sigma}}$ as
$\bm{\mathit{\Sigma}}=\begin{bmatrix}\bm{G}_{11}&\bm{{0}}_{3\times 3}&\bm{{0}}_{3\times 3}&\bm{G}_{12}&\bm{{0}}_{3\times 3}&\bm{{0}}_{3\times 3}\\ \bm{{0}}_{3\times 3}&\bm{G}_{11}&\bm{{0}}_{3\times 3}&\bm{{0}}_{3\times 3}&\bm{G}_{12}&\bm{{0}}_{3\times 3}\\ \bm{{0}}_{3\times 3}&\bm{{0}}_{3\times 3}&\bm{G}_{11}&\bm{{0}}_{3\times 3}&\bm{{0}}_{3\times 3}&\bm{G}_{12}\\ \bm{G}_{12}&\bm{{0}}_{3\times 3}&\bm{{0}}_{3\times 3}&\bm{G}_{22}&\bm{{0}}_{3\times 3}&\bm{{0}}_{3\times 3}\\ \bm{{0}}_{3\times 3}&\bm{G}_{12}&\bm{{0}}_{3\times 3}&\bm{{0}}_{3\times 3}&\bm{G}_{22}&\bm{{0}}_{3\times 3}\\ \bm{{0}}_{3\times 3}&\bm{{0}}_{3\times 3}&\bm{G}_{12}&\bm{{0}}_{3\times 3}&\bm{{0}}_{3\times 3}&\bm{G}_{22}\end{bmatrix},$
where under the proportional covariance, for $\bm{V}$ with Mat($\infty$) type structure, we have
$\bm{G}_{lm}=\begin{bmatrix}\gamma_{l,m}&\gamma_{l,m}r_{(1)}&\gamma_{l,m}r_{(1)}^{4}\\ \gamma_{l,m}r_{(1)}&\gamma_{l,m}&\gamma_{l,m}r_{(1)}\\ \gamma_{l,m}r_{(1)}^{4}&\gamma_{l,m}r_{(1)}&\gamma_{l,m}\end{bmatrix}.$
Here, $10<r_{(1)}<$ 1, $\gamma_{l,m}>0$, $1\leq l<m\leq 2$, and $\gamma_{11}\gamma_{22}>\gamma^{2}_{12}$. We can see that the matrices $\bm{G}_{11}$, $\bm{G}_{12}$ and $\bm{G}_{22}$ are proportional to each other. Under the generalized Markov-type covariance, for the NS1 structure of $\bm{\mathit{\Sigma}}$, we have
$\bm{G}_{11}=\sigma_{11}\begin{bmatrix}1&r_{(1)}&r_{(1)}^{2}\\ r_{(1)}&1&r_{(1)}\\ r_{(1)}^{2}&r_{(1)}&1\end{bmatrix},\leavevmode\nobreak\ \bm{G}_{12}=\rho\sqrt{\sigma_{11}\sigma_{22}}\begin{bmatrix}1&r_{(1)}&r_{(1)}^{2}\\ r_{(1)}&1&r_{(1)}\\ r_{(1)}^{2}&r_{(1)}&1\end{bmatrix}\text{ and}\\ \bm{G}_{22}=\sigma_{22}\begin{bmatrix}1&r_{(1)}\left(\rho^{2}+r_{(1)}\left(1-\rho^{2}\right)\right)&r_{(1)}^{2}\left(\rho^{2}+r^{2}_{(1)}\left(1-\rho^{2}\right)\right)\\ r_{(1)}\left(\rho^{2}+r_{(1)}\left(1-\rho^{2}\right)\right)&1&r_{(1)}\left(\rho^{2}+r_{(1)}\left(1-\rho^{2}\right)\right)\\ r_{(1)}^{2}\left(\rho^{2}+r^{2}_{(1)}\left(1-\rho^{2}\right)\right)&r_{(1)}\left(\rho^{2}+r_{(1)}\left(1-\rho^{2}\right)\right)&1\end{bmatrix},$
where $\sigma_{11},\sigma_{22}>0$, $10<|\rho|<$ 1 and $10<r_{(1)}<$ 1. It can be easily observed that the matrices $\bm{G}_{11}$ and $\bm{G}_{12}$ are proportional to each other, but for $r_{(1)}\neq 0$, the matrix $\bm{G}_{22}$ is not proportional to $\bm{G}_{11}$.
The choices presented in Table 1 and Table 2 are not exhaustive. We could have chosen any matrix for $\bm{V}$ (proportional case) as long as it is positive definite and symmetric, while in the Markovian case, $\bm{V}_{1}$ needs to be positive definite and symmetric while $\bm{V}_{R}$ can be any valid correlation matrix.
@@ 3 Search for efficient designs
We next determine a trace optimal/efficient design for the direct effects in model (2.2). Using the idea as provided by Bate and Jones, (2006), we say that a design $d^{*}\in\mathcal{D}$, where $\mathcal{D}$ is a subclass of designs, is a trace optimal/efficient design for the parameters of interest over $\mathcal{D}$ if $d^{*}$ maximizes the trace of the corresponding information matrix over $\mathcal{D}$.
For any design $d\in\Omega_{t,n,p}$, let $\bm{C}_{d(s1)}$ and $\bm{C}_{d(s2)}$ represent the information matrices of the direct effects for the proportional and generalized Markov-type structures, respectively. Lemma S.1 and Lemma S.3 show that the two information matrices have the following representations:
@@
$\bm{C}^{{}^{\prime}}_{{d(s2)}(21)(1)}=\left(\bm{I}_{2}\otimes\bm{T}_{d}^{{}^{\prime}}\right)\bm{A}^{*}\left(\bm{I}_{2}\otimes\bm{F}_{d}\bm{H}_{t}\right)$, $\bm{C}_{{d(s2)}(22)(1)}=\left(\bm{I}_{2}\otimes\bm{F}^{{}^{\prime}}_{d}\bm{H}_{t}\right)\bm{A}^{*}\left(\bm{I}_{2}\otimes\bm{F}_{d}\bm{H}_{t}\right)$, and the matrix $\bm{A}^{*}$ is $\bm{\mathit{\Sigma}}^{-1/2}pr^{\perp}\left(\bm{\mathit{\Sigma}}^{-1/2}\bm{Z_{1}}\right)\bm{\mathit{\Sigma}}^{-1/2}$. Here, $\bm{H}_{t}=\bm{I}_{t}-\frac{1}{t}\bm{{1}}_{t}\bm{{1}}{t}^{{}^{\prime}}$. Note that $\bm{A}^{*}$ depends on the structure of $\bm{\mathit{\Sigma}}$ and $\bm{C}_{d(s1)}$ is a $gt\times gt$ matrix, while $\bm{C}_{d(s2)}$ is a $2t\times 2t$ matrix.
@@ 3.1 Proportional covariance structure
$\bm{A}^{*}$ $=\bm{\mathit{\Gamma}}^{-1}\otimes\left(\bm{H}_{n}\otimes\bm{V}^{*}\right),$
@@
$\bm{C}_{d(uni)(12)}=\bm{C}_{d(uni)(21)}^{{}^{\prime}}=\bm{T}_{d}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{V}^{*}\right)\bm{F}_{d}$, and $\bm{C}_{d(uni)(22)}=\bm{F}_{d}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{V}^{*}\right)\bm{F}_{d}$. Here $\bm{V}^{*}=\bm{V}^{-1}-\left(\bm{{1}}_{p}^{{}^{\prime}}\bm{V}^{-1}\bm{{1}}_{p}\right)^{-1}\bm{V}^{-1}\bm{J}_{p\times p}\bm{V}^{-1}$.
@@
It should be noted that $\bm{C}_{d(uni)}$ is the information matrix for the direct effects when we have a univariate version of the model (2.1), i.e., $g=1$. To check if $OA_{I}\left(n=\lambda t\left(t-1\right),p=t,\right.$
$\left.t,2\right)$, where $\lambda$ is a positive integer and $t\geq 3$, is universally optimal, we explore complete symmetricity of the corresponding information matrix for $g>1$ case.
@@
where $\gamma^{(kk^{\prime})}$ is the $(k,k^{\prime})^{th}$ element of the matrix $\bm{\mathit{\Gamma}}^{-1}$. However, this condition is also seen to be violated for the $g>1$ setup (see Proof of Remark 3.1). Note, for a multivariate setup with no correlation between responses and a much simpler covariance structure, Niphadkar and Mukhopadhyay, (2023) have successfully obtained universally optimal designs. However, this is not true when the responses are correlated. Hence for the multivariate response case, we resort to identifying an efficient design in the next theorem.
@@
$\mathcal{D}^{(1)}_{t,n=\lambda t(t-1),p=t}$ is a class of binary designs with $p=t$, $\lambda$ is a positive integer and $t\geq 3$. It has been shown by Kunert and Martin, (2000) that $d^{*}$ is universally optimal for the direct effects over $\mathcal{D}^{(1)}_{t,n=\lambda t(t-1),p=t}$ in the univariate response case. Note that the universal optimality of a design over a subclass of designs implies that the design maximizes the trace of the corresponding information matrix over the same subclass of designs. Hence, $d^{*}$ maximizes $tr\left(\bm{C}_{d(uni)}\right)$ over $\mathcal{D}^{(1)}_{t,n=\lambda t(t-1),p=t}$, where $\bm{C}_{d(uni)}$ is the information matrix of the direct effects for the univariate response case. For $g>1$, using the expression of $\bm{C}_{d(s1)}$ from Theorem 3.1, we get
$tr\left(\bm{C}_{d(s1)}\right)$ $=tr\left(\bm{\mathit{\Gamma}}^{-1}\right)tr\left(\bm{C}_{d(uni)}\right).$ (3.4)
Here, $\bm{\mathit{\Gamma}}^{-1}$ is a positive definite matrix and thus $tr\left(\bm{\mathit{\Gamma}}^{-1}\right)>0$. So from (3.4), we get that for $g>1$, $d^{*}$ maximizes $tr\left(\bm{C}_{d(s1)}\right)$ over $\mathcal{D}^{(1)}_{t,n=\lambda t(t-1),p=t}$. Thus $d^{*}$ is trace optimal/efficient for the direct effects over $\mathcal{D}^{(1)}_{t,n=\lambda t(t-1),p=t}$ for the $g>1$ case. ∎
@@ 3.2 Generalized Markov-type covariance structure
For the Markovian structure, we are able to show that the $OA_{I}\left(n=\lambda t\left(t-1\right),p=t,t,2\right)$, where $\lambda$ is a positive integer and $t\geq 3$, is a highly efficient design. To evaluate the trace optimality/efficiency of these $OA_{I}\left(n=\lambda t\left(t-1\right),p=t,t,2\right)$ designs, we obtain an upper bound of $tr\left(\bm{C}_{d(s2)}\right)$, and state a necessary and sufficient condition for a $OA_{I}\left(n=\lambda t\left(t-1\right),p=t,t,2\right)$, to attain this upper bound of $tr\left(\bm{C}_{d(s2)}\right)$. Note, that if this upper bound is attained then the design maximizes $tr\left(\bm{C}_{d(s2)}\right)$ and hence is an efficient design. Also, for various choices of $\bm{V}_{1}$, $\bm{V}_{R}$ and $\rho$, we study the trace of the information matrix for the $OA_{I}\left(n=\lambda t\left(t-1\right),p=t,t,2\right)$ and the derived upper bound.
Proceeding as before, we see that the process of obtaining $\bm{A}^{*}$ is more complicated as compared to the proportional structure. This is mainly because under the Markovian structure, $\bm{\mathit{\Sigma}}_{22}$ is not proportional to $\bm{\mathit{\Sigma}}_{11}$. By laborious calculation, (see Lemma S.4), we obtain the expression of the matrix $\bm{A}^{*}$ under the generalized Markov-type structure as
$\bm{A}^{*}$ $=\begin{bmatrix}\bm{H}_{n}\otimes\bm{\mathit{\Omega}}_{1}&-\bm{H}_{n}\otimes\bm{\mathit{\Omega}}_{2}\\ -\bm{H}_{n}\otimes\bm{\mathit{\Omega}}_{2}&\bm{H}_{n}\otimes\bm{\mathit{\Omega}}_{4}\end{bmatrix},$
where $\bm{\mathit{\Omega}}_{1}=\bm{V}_{1}^{*}+\frac{\bar{\rho}^{2}}{\sigma_{12}}\bm{V}_{R}^{*}$, $\bm{\mathit{\Omega}}_{2}=\frac{\bar{\rho}}{\sigma_{12}}\bm{V}_{R}^{*}$ and $\bm{\mathit{\Omega}}_{4}=\frac{1}{\sigma_{12}}\bm{V}_{R}^{*}$. Here, $\bm{V}_{1}^{*}=\bm{V}_{1}^{-1}-\left(\bm{{1}}_{p}^{{}^{\prime}}\bm{V}_{1}^{-1}\bm{{1}}_{p}\right)^{-1}\bm{V}_{1}^{-1}\times$
$\bm{J}_{p\times p}\bm{V}_{1}^{-1}$, $\bm{V}_{R}^{*}=\bm{V}_{R}^{-1}-\left(\bm{{1}}_{p}^{{}^{\prime}}\bm{V}_{R}^{-1}\bm{{1}}_{p}\right)^{-1}\bm{V}_{R}^{-1}\bm{J}_{p\times p}\bm{V}_{R}^{-1}$, $\bm{H}_{n}=\bm{I}_{n}-\frac{1}{n}\bm{{1}}_{n}\bm{{1}}_{n}^{{}^{\prime}}$, $\bm{J}_{p\times p}=\bm{{1}}_{p}\bm{{1}}_{p}^{{}^{\prime}}$, $\bar{\rho}=\rho\sqrt{\frac{\sigma_{22}}{\sigma_{11}}}$ and $\sigma_{12}=\sigma_{22}\left(1-\rho^{2}\right)$.